# In Situ Training of Implicit Neural Compressors for Scientific Simulations via Sketch-Based Regularization

### v2

@@ 1 Introduction
Neural network-based compression methods can capture non-linear phenomena and sharp features, such as physical discontinuities, that pose difficulties for traditional linear dimensionality reduction techniques, such as those based on singular value decomposition (SVD). Several neural compression approaches exist, which can compress the data directly or learn a reduced dimensionality representation, but all require optimizing neural network weights under some training strategy. One such method, implicit neural representation (INR) [park2019deepsdf, genova2019learning, sitzmann2020implicit], is an implicit, continuous, and differentiable neural network model that, conveniently, only requires access to the space-time coordinates of points, circumventing the connectivity structure entirely. INR, with some specific recent developments, has shown outstanding accuracy in representing a diverse set of signals; see, e.g., [sitzmann2020implicit, zhang2021implicit, pan2023neural, han2023kd, sales2024implicit] or Figure˜2 wherein we illustrate sample compression results for INRs trained offline on two complex simulation datasets.
@@
Available data samples are stored in what we will refer to as the buffer, which in the extreme case is only large enough to store a single sample, though we will typically target the case when the buffer can hold the equivalent of a few dozen data samples. Altering the size of this buffer, along with varying how many samples arrive together or what information from the samples is stored, allows one to interpolate between the pure online and offline settings. Our notion of samples is general and need not be a single number; in a typical example, a data sample is the entire spatial field at a given time (a snapshot), arising from a time-dependent partial differential equation (PDE) simulation. Non-stationary training for traditional machine learning techniques has been widely studied, but often less so in the case of neural networks, where it is typically referred to as continual learning. Hypernetworks, [ha2017hypernetworks], which are neural network architectures that learn to generate weights for a target network, are perhaps the most successful and well-known method for continual learning. Appropriated from multi-task learning [caruana1997multitask], they allow for a natural extension to in situ learning by equating one or multiple time steps with a task.
@@ 1.1 Preliminaries and Related Work
Although our core methods are based on points $(\bm{x},t)\in\mathbb{R}^{d+1}$, for training, we will often consider a single sample to be a snapshot, i.e., $\bm{U}_{t}=\bm{u}(\bm{X},t)$, $\bm{U}_{t}\in\mathbb{R}^{n\times c}$. We then define the reconstruction of this snapshot, via a neural network $\bm{\Phi}:\mathbb{R}^{d+1}\to\mathbb{R}^{c}$ parameterized by $\bm{\theta}$, as $\bm{\widetilde{U}}_{t}:=\bm{\Phi}(\bm{X},t;\bm{\theta})$. The concatenation of these into the full $T\times n\times c$ dateset and its reconstruction are given as $\bm{U}=[\bm{U}_{1};\cdots;\bm{U}_{T}]$ and $\bm{\widetilde{U}}=[\bm{\widetilde{U}}_{1};\cdots;\bm{\widetilde{U}}_{T}]$, respectively.
@@
Note that this does not include the required storage for the mesh or network definition, or intermediate storage for the in situ learning scheme. The formula assumes consistent floating point representation of all numbers, but should be adjusted in the obvious way if, for example, some of the network weights are represented in different precisions.
@@
INR is a straightforward form of supervised deep learning to train a neural network $\bm{\Phi}$ to approximate a target signal $\bm{f}$, and is the building block framework upon which have been built popular extensions, such as PINNs [raissi2019physics]. The key concept is to build a neural network that views input as discrete samples of a continuous object, so that dimensionality does not increase with resolution. Typically, the network in question is of a simple, densely-connected design, and the signal is a function of spatial or spatio-temporal coordinates (for notational simplicity, in this subsection we assume a purely spatial domain and hence denote the coordinates by $\bm{x}$). To be more precise, we consider a signal $\bm{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ that maps coordinates in $d$ dimensions to a vector of $c$ features. Examples include 2D image pixel coordinates mapping to RGB colors, or mesh coordinates and time mapping to the flow field of a numerical PDE solution. Section˜1.1 defines the simplest INR with $L$ feedforward layers and non-linear activation function $\sigma$, but one should note that more complicated architectural features, such as residual connections, can be included:
@@
The activation function $\sigma$ has been shown to play a significant role in the performance of INR. Most notably, the work of [sitzmann2020implicit] demonstrated that a sinusoidal activation with a specific random initialization of the network weights is a vastly superior option compared to ReLU, Tanh, or other popular alternatives, calling their framework a SIREN network. Extensions of the periodic activation function have been considered by, e.g., [saragadam2023wire, roddenberry2023implicit], which introduce and analyze complex wavelets. See the comprehensive review [essakine2024we] for further details on the current state of INR.
@@
Hypernetworks are a simple but powerful concept in deep learning wherein one network, the hypernetwork, is used to generate weights for a target network. Originally developed by [ha2017hypernetworks] as a form of weight sharing for convolutional and recurrent models, they have since been integrated with a variety of deep learning pipelines. Typically, the hypernetwork is much smaller than the target network and maps from an encoding of a specific task; for example, language pairs in translation. Less complicated architectures map directly to the full vector of target weights, but for large target models, this can be inefficient. Other approaches generate weights in chunks or via scaling factors. For details on the extensive work surrounding hypernetworks, see the review [chauhan2024brief].
Of particular relevance to our work is that of [pan2023neural], which combines hypernetworks and INR for learning low-dimensional representations of large scientific datasets. We use a similar sine-based INR with residual connections and generate time-dependent weights in the hypernetwork. Compression is not the focus of their work, and all models are trained offline, unlike our focus on in situ training and compression. Further, many of the stated disadvantages in the aforementioned work stem from a comparison to the SVD, which does not require hyperparameter tuning or network optimization. However, when the signal of interest exhibits a large Kolmogorov $n$-width, as in compressible and advection-dominated flows, linear dimensionality reduction techniques, such as the SVD, cease to be effective, thus motivating the use of non-linear dimensionality reduction, e.g., methods based on neural networks [lee2020model]. We also find that most of the remaining disadvantages, such as improving parameter efficiency or reducing training time, are deployment tasks, which can be tackled in isolation from the core compression problem.
@@
The machine learning community has used many synonymous or related terms, such as online learning, incremental learning, and continual learning, to describe training on data with a non-stationary distribution. This setting stands in contrast to the usual offline regime of most deep learning, where one assumes unfettered access to a fixed training dataset, which is used to optimize a model that is then deployed or tested on other data. In reality, a spectrum exists between fully offline learning and fully online learning, where samples arrive once and one at a time. Possible variations include multiple samples available at one time or the ability to store some data offline. The key difficulty lies in avoiding catastrophic forgetting, i.e., maintaining good performance for samples seen earlier in the training process [kirkpatrick2017overcoming].
Hypernetworks are especially adept at tackling these non-stationary problems, as the work of [von2019continual] highlights. To address catastrophic forgetting, they introduce a regularization term that seeks to maintain the hypernetwork output on previously learned tasks while learning a new task. A copy of the model is saved after learning each new task, which is used to reproduce the output of the hypernetwork on previous tasks. This method yields exceptional performance and limited forgetting, but the maximum number of tasks never exceeds 100. In our context, a task is a time step, and 100 is far fewer than the number of time steps often encountered in PDE simulations.
Our regularization approach using sketched spatial snapshots is more similar to the experience replay or rehearsal techniques commonly seen in reinforcement learning [rolnick2019experience] and throughout continual learning [wang2024comprehensive], though the technical details are notably different.
@@
A large portion of this effort has been focused on error-bounded schemes for highly accurate compression of structured data, but at relatively small compression rates. Well-known tools include ZFP [lindstrom2014zfp] and SZ3 [liang2023sz], along with the recently introduced MFZ [doherty2024mfz] that extends to non-uniform data. See the recent review [di2025survey] for further details on error-bounded methods.
@@
As mentioned earlier, significant effort has been put towards linear low-rank approximation methods for compression, motivated by the best [^2] In any unitarily invariant norm. rank-$k$ approximation from the SVD [eckart1936approximation]. Deterministic methods are often too computationally expensive for large datasets, so many methods turn to randomized decompositions. This employs randomized sketching [halko2011finding] — similar to what we consider in this work — but directly for compression as opposed to indirectly through regularization. We highlight the single-pass randomized methods using the SVD [yu2017single] and interpolative decomposition (ID) [dunton2020pass, pacella2022task, li2025online] for their focus on in situ compression.
@@
One of the more significant directions of neural compression research uses INR in implicit neural compression (INC). Here, compression is accomplished by simply requiring that the size of the network is smaller than the size of the data so that Eq.˜1 is greater than $1$. Other approaches to compression include autoencoders, with recent developments to enable arbitrarily structured meshes [doherty2024quadconv]. A mixed autoencoder INR approach is also developed in [pham2023autoencoding]. The work of [yang2023tinc] introduces a blocking tree-based method with parameter sharing, and multiscale representations have also been considered in [saragadam2022miner, tang2024ecnr].
@@
Most INC research has been focused on images [strumpler2022implicit, dupont2021coin] or video [zhang2021implicit, chen2021nerv], and some for general signals [pistilli2022signal]. More relevant to our task are [lu2021compressive, sales2024implicit], which apply INC to three and four-dimensional volumetric data from scientific simulations. Importantly, these works consider the offline scenario, incurring large training times and often using the entire dataset for meta-learning in a pre-training stage. Most also incorporate some form of quantization to reduce the model parameters, and thus increase the compression rate further.
@@
A hybrid approach is considered in [madireddy2021], where a convolutional neural network is used to correct JPEG-based compression via artifact removal. A pre-trained model is fine-tuned in situ, in part using a buffer of stored data, before it is deployed to compress, i.e., remove artifacts, in a full simulation. While their use of experience replay ideas is similar to ours, the fundamental process is significantly different, given our sketch-based approach. Additionally, similar to many offline neural compression methods, their work requires significant processing outside of the in situ learning stage.
@@
The most similar work to our own is the knowledge distillation approach of [han2023kd] (KD-INR), which also uses INR, but not hypernetworks. The difficulty associated with in situ learning is avoided by breaking the problem down into individual offline learning tasks. In particular, a new time-independent INR is trained at every time step, and then these are all distilled into a single time-dependent INR using a final knowledge transfer training stage. This final stage requires an extra training step, relative to our method, after the simulation has concluded. Additionally, this approach uses an offline pre-processing step to normalize the data and does not consider multi-channel training, although it is unclear how this impacts performance. Without the offline normalization step, KD-INR is also an in situ method, but it incurs different costs from our own approach, due to saving different objects. We did not compare directly to KD-INR because the code and data were not publicly available, so this is left to future work.
@@ 2 In Situ Training
To ensure that the data is used most effectively in this single-pass regime, we first assume that we have some amount of available working memory, large enough to hold multiple snapshots at once. This memory is used to implement a buffer which holds data extracted from the simulation (Fig.˜3). The first portion of this buffer holds full snapshots in a queue of size $T_{f}\ll T$ — in fact, our numerical experiments use $T_{f}=1$. Newly generated snapshots are pushed into the queue, replacing old ones as the maximum size is reached. The second portion of the buffer contains sketched snapshots; the overall size of this buffer is not large, but its length $T_{s}\leq T$ can be quite long since the dimension of each sketch is very small. As we shall explain in further in Section˜2.1, sketching refers to a linear dimensionality reduction technique wherein a linear operator $\bm{S}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ projects an $n$-dimensional vector onto a $k$-dimensional subspace, with $k\ll n$, such that some information in the original data is preserved. The total size of the stored buffer data is then given as follows:
@@
where $\lambda>0$ weights the sketch regularization term. In the case of $T_{f}=1$ the first term simplifies significantly. The mini-batch samples are always drawn anew at every iteration of the optimization algorithm, but the sketch $\bm{S}_{t^{\prime}}$ is only drawn once per time-step and then reused. It is also important to note that the product $\bm{S}_{t_{m}}\bm{U}_{t_{m}}$ is stored in the sketched buffer, as it cannot be reformed since we may not have saved the full snapshot. The sketched reconstruction $\bm{S}_{t_{m}}\bm{\widetilde{U}}_{t_{m}}$ is formed by sketching the full INR reconstruction, because only subsampling sketches can be applied by evaluating the network at a specific set (the subsample) of points.
@@
$\bm{\Phi}\leftarrow\text{build\_model}(~)$ ;
@@
$\bm{B}_{f}\leftarrow\text{queue}(T_{f})$ ;
@@
$\bm{B}_{s}\leftarrow\text{queue}(T_{s})$ ;
@@
    $\bm{S}_{t}\leftarrow\text{construct\_sketch}(k)$ ;
    // Construct sketch
    $\text{push}(\bm{B}_{f},\bm{U}_{t})$ ;
@@
    $\bm{U}_{f}\leftarrow\text{minibatch}(\bm{B}_{f},b_{f})$ ;
@@
    $\bm{U}_{s}\leftarrow\text{minibatch}(\bm{B}_{s},b_{s})$ ;
@@
    $\text{update\_model}(\bm{\Phi};\bm{U}_{f},\bm{U}_{s})$ ;

### v3

@@ 1 Introduction
Neural network-based compression methods can capture non-linear phenomena and sharp features, such as physical discontinuities, that pose difficulties for traditional linear dimensionality reduction techniques, such as those based on singular value decomposition (SVD). Several neural compression approaches exist, which can compress the data directly or learn a reduced dimensionality representation, but all require optimizing neural network weights under some training strategy. One such method, implicit neural representation (INR) [40, 20, 50], is an implicit, continuous, and differentiable neural network model that, conveniently, only requires access to the space-time coordinates of points, circumventing the connectivity structure entirely. INR, with some specific recent developments, has shown outstanding accuracy in representing a diverse set of signals; see, e.g., [50, 58, 39, 24, 46] or Figure˜2 wherein we illustrate sample compression results for INRs trained offline on two complex simulation datasets.
@@
Available data samples are stored in what we will refer to as the buffer, which in the extreme case is only large enough to store a single sample, though we will typically target the case when the buffer can hold the equivalent of a few dozen data samples. Altering the size of this buffer, along with varying how many samples arrive together or what information from the samples is stored, allows one to interpolate between the pure online and offline settings. Our notion of samples is general and need not be a single number; in a typical example, a data sample is the entire spatial field at a given time (a snapshot), arising from a time-dependent partial differential equation (PDE) simulation. Non-stationary training for traditional machine learning techniques has been widely studied, but often less so in the case of neural networks, where it is typically referred to as continual learning. Hypernetworks, [22], which are neural network architectures that learn to generate weights for a target network, are perhaps the most successful and well-known method for continual learning. Appropriated from multi-task learning [6], they allow for a natural extension to in situ learning by equating one or multiple time steps with a task.
@@ 1.1 Preliminaries and Related Work
Although our core methods are based on points $(\bm{x},t)\in\mathbb{R}^{d+1}$, for training, we will often consider a single sample to be a snapshot, i.e., $\bm{U}_{t}:=\bm{u}(\bm{X},t)$, $\bm{U}_{t}\in\mathbb{R}^{n\times c}$. We then define the reconstruction of this snapshot, via a neural network $\bm{\Phi}:\mathbb{R}^{d+1}\to\mathbb{R}^{c}$ parameterized by $\bm{\theta}$, as $\bm{\widetilde{U}}_{t}:=\bm{\Phi}(\bm{X},t;\bm{\theta})$. The concatenation of these into the full $T\times n\times c$ dateset and its reconstruction are given as $\bm{U}=[\bm{U}_{1};\cdots;\bm{U}_{T}]$ and $\bm{\widetilde{U}}=[\bm{\widetilde{U}}_{1};\cdots;\bm{\widetilde{U}}_{T}]$, respectively. In certain contexts, we will drop the dependence of $\bm{\Phi}$ on one or more of its arguments to make clear that these are fixed. For example, we write $\bm{\Phi}(\bm{\theta})$ when considering the network as a function of its parameters for fixed input coordinates.
@@
Note that this does not include the required storage for the mesh or network definition, or intermediate storage for the in situ learning scheme. The formula assumes consistent floating-point representation of all numbers, but should be adjusted in the obvious way if, for example, some of the network weights are represented in different precisions.
@@
INR is a straightforward form of supervised deep learning to train a neural network $\bm{\Phi}$ to approximate a target signal $\bm{f}$, and is the building block framework upon which popular extensions, such as PINNs [43], have been built. The key concept is to build a neural network that views input as discrete samples of a continuous object, so that dimensionality does not increase with resolution. Typically, the network in question is of a simple, densely-connected design, and the signal is a function of spatial or spatio-temporal coordinates (for notational simplicity, in this subsection we assume a purely spatial domain and hence denote the coordinates by $\bm{x}$). To be more precise, we consider a signal $\bm{f}:\mathbb{R}^{d}\to\mathbb{R}^{c}$ that maps coordinates in $d$ dimensions to a vector of $c$ features. Examples include 2D image pixel coordinates mapping to RGB colors, or mesh coordinates and time mapping to the flow field of a numerical PDE solution. Section˜1.1 defines the simplest INR with $L$ feedforward layers and non-linear activation function $\sigma$, but one should note that more complicated architectural features, such as residual connections, can be included:
@@
The activation function $\sigma$ has been shown to play a significant role in the performance of INR. Most notably, the work of [50] demonstrated that a sinusoidal activation with a specific random initialization of the network weights is a vastly superior option compared to ReLU, Tanh, or other popular alternatives, calling their framework a SIREN network. Extensions of the periodic activation function have been considered by, e.g., [47, 44], which introduce and analyze complex wavelets. See the comprehensive review [18] for further details on the current state of INR.
@@
Hypernetworks are a simple but powerful concept in deep learning wherein one network, the hypernetwork, is used to generate weights for a target network. Originally developed by [22] as a form of weight sharing for convolutional and recurrent models, they have since been integrated with a variety of deep learning pipelines. Often, the hypernetwork is much smaller than the target network [^2] Although this is heavily dependent on the mapping and is not the case in our models. and maps from an encoding of a specific task; for example, language pairs in translation. Less complicated architectures map directly to the full vector of target weights, but for large target models, this can be inefficient. Other approaches generate weights in chunks or via scaling factors. For details on the extensive work surrounding hypernetworks, see the review [7].
Of particular relevance to our work is that of [39], which combines hypernetworks and INR for learning low-dimensional representations of large scientific datasets. We use a similar sine-based INR with residual connections and generate time-dependent weights in the hypernetwork. Compression is not the focus of their work, and all models are trained offline, unlike our focus on in situ training and compression. Further, many of the stated disadvantages in the aforementioned work stem from a comparison to the SVD, which does not require hyperparameter tuning or network optimization. However, when the signal of interest exhibits a large Kolmogorov $n$-width, as in compressible and advection-dominated flows, linear dimensionality reduction techniques, such as the SVD, cease to be effective, thus motivating the use of non-linear dimensionality reduction, e.g., methods based on neural networks [30]. We also find that most of the remaining disadvantages, such as improving parameter efficiency or reducing training time, are deployment tasks, which can be tackled in isolation from the core compression problem.
@@
The machine learning community has used many synonymous or related terms, such as online learning, incremental learning, and continual learning, to describe training on data with a non-stationary distribution. This setting stands in contrast to the usual offline regime of most deep learning, where one assumes unfettered access to a fixed training dataset, which is used to optimize a model that is then deployed or tested on other data. In reality, a spectrum exists between fully offline learning and fully online learning, where samples arrive once and one at a time. Possible variations include multiple samples available at one time or the ability to store some data offline. The key difficulty lies in avoiding catastrophic forgetting, i.e., maintaining good performance for samples seen earlier in the training process [28].
Hypernetworks are especially adept at tackling these non-stationary problems, as the work of [53] highlights. To address catastrophic forgetting, they introduce a regularization term that seeks to maintain the hypernetwork output on previously learned tasks while learning a new task. A copy of the model is saved after learning each new task, which is used to reproduce the output of the hypernetwork on previous tasks. This method yields exceptional performance and limited forgetting, but the maximum number of tasks never exceeds 100. In our context, a task is a time step, and 100 is far fewer than the number of time steps often encountered in PDE simulations.
Our regularization approach using sketched spatial snapshots is more similar to the experience replay or rehearsal techniques commonly seen in reinforcement learning [45] and throughout continual learning [54], though the technical details are notably different.
@@
A large portion of this effort has been focused on error-bounded schemes for highly accurate compression of structured data, but at relatively small compression rates. Well-known tools include ZFP [35] and SZ3 [34], along with the recently introduced MFZ [11] that extends to non-uniform data. See the recent review [9] for further details on error-bounded methods.
@@
As mentioned earlier, significant effort has been put towards linear low-rank approximation methods for compression, motivated by the best [^3] In any unitarily invariant norm. rank-$k$ approximation from the SVD [16]. Deterministic methods are often too computationally expensive for large datasets, so many methods turn to randomized decompositions. This employs randomized sketching [23] — similar to what we consider in this work — but directly for compression as opposed to indirectly through regularization. We highlight the single-pass randomized methods using the SVD [57] and interpolative decomposition (ID) [13, 38, 31] for their focus on in situ compression.
@@
One of the more significant directions of neural compression research uses INR in implicit neural compression (INC). Here, compression is accomplished by simply requiring that the size of the network is smaller than the size of the data so that Eq.˜1 is greater than $1$. Other approaches to compression include autoencoders, with recent developments to enable arbitrarily structured meshes [12]. A mixed autoencoder INR approach is also developed in [41]. The work of [56] introduces a blocking tree-based method with parameter sharing, and multiscale representations have also been considered in [48, 52].
@@
Most INC research has been focused on images [51, 15] or video [58, 8], and some for general signals [42]. More relevant to our task are [36, 46], which apply INC to three and four-dimensional volumetric data from scientific simulations. Importantly, these works consider the offline scenario, incurring large training times and often using the entire dataset for meta-learning in a pre-training stage. Most also incorporate some form of quantization to reduce the model parameters, and thus increase the compression rate further.
@@
A hybrid approach is considered in [37], where a convolutional neural network is used to correct JPEG-based compression via artifact removal. A pre-trained model is fine-tuned in situ, in part using a buffer of stored data, before it is deployed to compress, i.e., remove artifacts, in a full simulation. While their use of experience replay ideas is similar to ours, the fundamental process is significantly different, given our sketch-based approach. Additionally, similar to many offline neural compression methods, their work requires significant processing outside of the in situ learning stage.
@@
The most similar work to our own is the knowledge distillation approach of [24] (KD-INR), which also uses INR, but not hypernetworks. The difficulty associated with in situ learning is avoided by breaking the problem down into individual offline learning tasks. In particular, a new time-independent INR is trained at every time step, and then these are all distilled into a single time-dependent INR using a final knowledge transfer training stage. This final stage requires an extra training step, relative to our method, after the simulation has concluded. Additionally, this approach uses an offline pre-processing step to normalize the data and does not consider multi-channel training, although it is unclear how this impacts performance. Without the offline normalization step, KD-INR is also an in situ method, but it incurs different costs from our own approach, due to saving different objects. We did not compare directly to KD-INR because the code and data were not publicly available, so this is left to future work.
@@ 2 In Situ Training
To ensure that the data is used most effectively in this single-pass regime, we first assume that we have some amount of available working memory, large enough to hold multiple snapshots at once. This memory is used to implement a buffer which holds data extracted from the simulation, visualized in Fig.˜3. The first portion of this buffer holds full snapshots in a queue of size $T_{f}\ll T$ — in fact, our numerical experiments use $T_{f}=1$. Newly generated snapshots are pushed into the queue, replacing old ones as the maximum size is reached. The second portion of the buffer contains sketched snapshots; the overall size of this buffer is not large, but its length $T_{s}\leq T$ can be quite long since the dimension of each sketch is very small. As we shall explain in further in Section˜2.1, sketching refers to a linear dimensionality reduction technique wherein a linear operator $\bm{S}:\mathbb{R}^{n\times c}\to\mathbb{R}^{k\times c}$ projects an $n$-dimensional vector onto a $k$-dimensional subspace, with $k\ll n$, such that some information in the original data is preserved. The total size of the stored buffer data is then given as follows:
@@
where $\lambda>0$ weights the sketch regularization term. In the case of $T_{f}=1$, the first term simplifies significantly. The mini-batch samples are always drawn anew at every iteration of the optimization algorithm, but the sketch $\bm{S}_{t^{\prime}}$ is only drawn once per time-step and then reused. It is also important to note that the product $\bm{S}_{t_{m}}\bm{U}_{t_{m}}$ is stored in the sketched buffer, as it cannot be reformed since we may not have saved the full snapshot. The sketched reconstruction $\bm{S}_{t_{m}}\bm{\widetilde{U}}_{t_{m}}$ is formed by sketching the full INR reconstruction, because only subsampling sketches can be applied by evaluating the network at a specific set (the subsample) of points.
@@
$\bm{\Phi}\leftarrow\text{build\_model}(~)$
@@
$\bm{B}_{f}\leftarrow\text{queue}(T_{f})$
@@
$\bm{B}_{s}\leftarrow\text{queue}(T_{s})$
@@
    $\bm{S}_{t}\leftarrow\text{construct\_sketch}(k)$
    // Construct sketch (storing seed)
    $\text{push}(\bm{B}_{f},\bm{U}_{t})$
@@
    $\bm{U}_{f}\leftarrow\text{minibatch}(\bm{B}_{f},b_{f})$
@@
     $\bm{\widehat{U}}_{s}\leftarrow\text{minibatch}(\bm{B}_{s},b_{s})$
@@
     $\text{compute\_loss}(\bm{\Phi};\bm{U}_{f},\bm{\widehat{U}}_{s})$
     // Via Eq.˜4 with explicit minibatches
     $\text{update\_model}(\bm{\Phi})$